### v2

# Supplementary Materials for: Cryogenic piezoelectric effects in thin film strontium titanate devices

Supplementary Materials for:
Cryogenic piezoelectric effects in thin film strontium titanate devices
Ahmed Khalil, et al
See pages - of Main.pdf
Supplementary Materials for
Cryogenic piezoelectric effects in thin film
strontium titanate devices
Ahmed Khalil et al.
Corresponding author. Email: Christian.Haffner@imec.be
This PDF file includes:
Supplementary Text
Figures S1 to S9
Tables S1 to S2
1 Piezoelectricity extraction
[…]
Fig. S1: IDT on SrTiO3 thin film. Microscope image of a representative IDT fabricated on a SrTiO3 thin film.
1.1 Electrical characterization and impedance extraction
All devices were characterized at a temperature of 5 K using an Attocube attoDRY800 system and Keysight (P5003B) vector network analyzer (VNA). Using the measured reflection coefficient $S_{11}(f)$, calibrated up to the device under test, we extract the IDT impedance as
$Z(f)=Z_{0}\frac{1+S_{11}(f)}{1-S_{11}(f)},$ (S1)
where $Z_{0}=50~\Omega$ is the characteristic impedance of the measurement system.
$R_{s}$$L_{s}$$C_{T}$$B_{m}$$G_{m}$$Y_{T}$
Fig. S2: Mason model and permittivity extraction. (a) Mason equivalent circuit model for the IDT electromechanical response. The typical measured resistance and wideband $\mathrm{Im}[Z]$ are shown in (b) and (c), respectively. (d) Shows the numerically simulated map from capacitance to SrTiO3 permittivity.
1.2 Mason model and electromechanical coupling
The electromechanical response of the IDTs is better understood using Mason’s equivalent circuit model, in which the total electrical admittance $Y(f)$ is expressed as the sum of a static capacitive contribution and a frequency-dependent motional response  (?):
$Y(f)=i\omega C_{T}+G_{m}(f)+iB_{m}(f),$ (S2)
where $C_{T}$ is the static (geometric) capacitance of the IDT, and $G_{m}(f)$ and $B_{m}(f)$ are the motional conductance and susceptance, respectively. The motional conductance $G_{m}(f)$ quantifies the conversion of electrical energy into mechanical (acoustic) energy and is given by
$G_{m}(f)=8K^{2}C_{T}Nf_{0}\left(\frac{\sin X}{X}\right)^{2}\equiv G_{0}\left(\frac{\sin X}{X}\right)^{2},\qquad X=N\pi\frac{|f-f_{0}|}{f_{0}},$ (S3)
where $K^{2}$ is the electromechanical coupling coefficient, $f_{0}$ is the acoustic resonance frequency, and $G_{0}$ denotes the peak motional conductance. By causality, the motional susceptance $B_{m}(f)$ is related to $G_{m}(f)$ via a Hilbert transform and is given by
$B_{m}(f)=G_{0}\frac{\sin(2X)-2X}{2X^{2}}.$ (S4)
Both $G_{m}(f)$ and $B_{m}(f)$ are narrow-band features centered around $f_{0}$ with a fractional bandwidth of $\Delta f/f_{0}\approx 0.885/N$. Fig. S2(a) shows a circuit representation of the Mason model with addition of parasitic resistance ($R_{s}$) and inductance ($L_{s}$). Fig. S2(b,c) show a typical measured resistance $\text{Re}[Z]$, and wide-band Im[Z][Z], respectively.
As evident from Fig. S2 (b), the parasitic resistance contributes a constant background to the measured response. In the following, we focus on extracting the motional resistance that appears on top of this background. At resonance, the motional resistance is given by
$R_{m}(f_{0})=\Re\{1/Y(f_{0})\}=\frac{G_{0}}{G_{0}^{2}+\left(\omega_{0}C_{T}\right)^{2}}.$ (S5)
The motional resistance is governed by the electromechanical coupling coefficient $K^{2}$, which itself is determined by the piezoelectric tensor ($\mathbf{e}$), permittivity ($\mathbf{\epsilon}$) and elasticity ($\mathbf{c}$) tensors of the SrTiO3 film. In bulk acoustic wave transducers, $K^{2}$ is given by $K^{2}=\frac{e_{33}^{2}}{c_{33}\varepsilon_{33}}$   (?). However, for surface acoustic wave (SAW) transducers, $K^{2}$ does not have a closed-form expression and can be regarded as a function $K^{2}=f(\mathbf{e,\epsilon,c})$. Therefore, we numerically simulate the IDT electromechanical response using 2D-FEM in COMSOL Multiphysics to model the Rayleigh wave and extract the SrTiO3 material properties. The entire number of fingers was simulated to calculate the motional resistance. Since the simulation is two dimensional, the shear waves (out-of-plane in the 2D simulation) is not captured. The required permittivity extraction is extracted using the COMSOL map in Fig. S2(d) as discussed earlier in the methods section. The elasticity tensor is taken from literature  (?, ?), namely, $c_{11}=337.23$ [GPa], $c_{12}=132.2$ [GPa], $c_{44}=115.34$ [GPa]. In the simulation, the elasticity of SrTiO3 was modeled as a ”cubic” tensor, where the diagonal elements satisfy $c_{11}=c_{22}=c_{33}$, the off-diagonal elements $c_{ij}$ (for $i\neq j$ and $i,j=1,2,3$) are equal, and the shear components satisfy $c_{44}=c_{55}=c_{66}$. This approach shows a good agreement with our acoustic dispersion results presented in the main text. The only free parameter is the piezoelectric tensor $\mathbf{e}$, which is adjusted to fit the amplitude of the measured motional resistance.
For Rayleigh-wave excitation, the dominant piezoelectric tensor components are $e_{33}$ and $e_{31}$, consistent with the predominantly in-plane displacement of Ti ions in the SrTiO3 thin films as discussed in section 1.4. The $e_{31}$ component is negative and can be physically understood as the transverse contraction of the unit cell accompanying expansion along the polar axis. For ABO3 perovskites, $|e_{31}|$ is typically at least one order of magnitude smaller than $e_{33}$  (?). Density functional theory calculations for SrTiO3 suggest even smaller values of $e_{31}$  (?). In our analysis, we therefore assume $e_{31}=-0.1\,e_{33}$, as an upper bound estimation for $e_{31}$ (lower bound for $e_{33}$) based on related perovskite materials. In summary, the extraction methodology for $e_{33,\mathrm{eff}}$ is as follows:
Analyzing wide-band data:
$\Im{[Z]}\xrightarrow{\text{ Fitting }}C[pF]\xrightarrow{\text{ COMSOL }}\epsilon_{\mathrm{STO}}$
Analyzing near-resonance data:
$\Re{[Z]}_{\text{No parasitics}}\xrightarrow{\operatorname{COMSOL}\left(\epsilon_{\mathrm{STO}}\right)}e_{33,\mathrm{eff}}$
1.3 $e_{15}$ extraction based on BAW
The transducer can also be considered as a lateral field excitation (LFE) transducer with parallel electrodes configuration. In this case, the analysis is simpler, the electromechanical coupling coefficient for thickness shear wave excitation is given by (?, ?)
$K_{15,\mathrm{eff}}=\frac{e_{15,\mathrm{eff}}}{\sqrt{c_{55}^{E}\varepsilon_{11}^{T}}}.$ (S6)
Additionally, the $K_{15,\mathrm{eff}}$ can be experimentally measured using the frequency separation between the parallel and series resonance frequencies
$K_{15,\mathrm{eff}}^{2}=\left(\frac{\pi}{2}\right)^{2}\frac{f_{p}-f_{s}}{f_{p}},$ (S7)
where $f_{p}$ is the parallel resonance frequency and $f_{s}$ is the series resonance frequency shown in Fig. S3.
[…]
The figure shows an FSR of 4.2 MHz due to the back-reflections from the Si substrate. In order to mitigate the uncertainty in calculating the frequency separation, the separation is extracted by averaging the upper and lower impedance envelopes. This procedure is justified by comparing the results of Mason’s model for bounded and unbounded transducers.
1.4 Acoustic wave excitation dependency on piezoelectric tensor
To understand the dependency of Rayleigh wave excitation on different piezoelectric tensor components, we solve the governing equations for acoustic wave as an eigenvalue problem to obtain the acoustic modes’ profiles and their corresponding eigenfrequencies (velocities) in COMSOL Multiphysics, namely (?),
$\nabla_{s}\left(\mathbf{c}\left[\nabla_{s}\right]^{T}\mathbf{u}+\mathbf{e}^{T}[\nabla]^{T}\phi\right)+\rho\omega^{2}\mathbf{u}$ $=0,$ (S8)
$\nabla\left(\mathbf{e}\left[\nabla_{s}\right]^{T}\mathbf{u}-\varepsilon[\nabla]^{T}\phi\right)$ $=0.$
where $\mathbf{u}$ is the displacement vector, $\phi$ is the electric potential, $\mathbf{c}$ is the elasticity tensor, $\mathbf{e}$ is the piezoelectric stress-tensor, $\boldsymbol{\varepsilon}$ is the permittivity tensor and $\rho$ is the density. For a given acoustic mode, the magnitude of the gap between dispersion curves calculated with and without piezoelectric coupling provides a direct measure of how efficiently that mode can be electrically excited. Piezoelectric coupling modifies the acoustic phase velocity; by reciprocity, this implies that an electric field applied along the corresponding crystallographic direction can efficiently drive the same acoustic mode. Guided by this principle, we perform single-domain simulations both including piezoelectric coupling, implemented via free-floating electric potential boundary conditions (BCs), and excluding piezoelectric coupling, enforced through grounded potential BCs. This procedure enables the systematic construction of slowness curves (inverse phase velocity) as a function of propagation direction and allows the contribution of individual piezoelectric tensor components to be isolated. Fig. S4a shows the slowness curves of the Rayleigh wave under free and grounded potential BCs, considering only the $e_{33}$ component. It follows that the electromechanical coupling can be calculated from the slowness curves as (?)
$K^{2}=2\times\left(1-\frac{v_{\mathrm{grounded}}}{v_{\mathrm{free}}}\right)\times 100\%.$ (S9)
Refer to caption(a)(b)
Fig. S4: Rayleigh wave slowness curves and coupling. (a) Slowness curves for Rayleigh wave with and without piezoelectric coupling accounting only $e_{33}$ component. (b) Rayleigh wave electromechanical coupling for different piezoelectric tensor elements.
Fig. S4(b) shows the dependency of the Rayleigh wave excitation on the piezoelectric coefficient $e_{33}$. The wave can be excited efficiently in the c-oriented domain but not in the a-oriented domain. It can be seen that the $e_{31}$ component slightly enhances the excitation, while the $e_{15}$ component doesn’t contribute to the Rayleigh wave excitation with 0 and 90 degrees IDTs. Similar interpretations can be drawn by merely looking at the rotated piezoelectric tensor. As a reminder, the piezoelectric tensor for tetragonal c-axis (Ti atom is out of the wafer’s plane) SrTiO3 is given by
$e]=\begin{bmatrix}0&0&0&0&e_{15}&0\\ 0&0&0&e_{15}&0&0\\ e_{31}&e_{31}&e_{33}&0&0&0\end{bmatrix}$. (S10)
For a c-axis film, the crystallographic axes $(1,2,3)$ are parallel to the electric-field coordinate system, corresponding to ($1=x$, $2=y$, and $3=z$). However, our grown SrTiO3 is a multidomain film with Ti atoms oriented in-plane (a-axis) either with 0 (regarded a-oriented) as
$e]_{a}^{0}=\begin{array}[]{cccccc}S_{1}\quad S_{2}\quad S_{3}\quad S_{4}\quad S_{5}\quad S_{6}\\ \left[\begin{array}[]{cccccc}0&0&0&0&0&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{15}\\ e_{31}&e_{33}&e_{31}&0&0&0\\ 0&0&0&e_{15}&0&0\\ \end{array}\right]\\ \end{array}\begin{gathered}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}E_{x}\\ E_{y}\\ E_{z}\end{gathered}$ (S11)
or with 90 (regarded c-oriented) as
$e]_{c}^{90}=\begin{array}[]{cccccc}S_{1}\quad S_{2}\quad S_{3}\quad S_{4}\quad S_{5}\quad S_{6}\\ \left[\begin{array}[]{cccccc}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{33}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{31}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{31}&0&0&0\\ 0&0&0&0&0&e_{15}\\ 0&0&0&0&e_{15}&0\\ \end{array}\right]\\ \end{array}\begin{gathered}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}E_{x}\\ E_{y}\\ E_{z}\end{gathered}$ (S12)
For a dominantly x-oriented electric field across $e]_{a}^{0}$ the $e_{15}$ now is at 1,6 position which couples x-field to shear x-y stress, consistent with the mode profile shown in Fig. S5a. Similarly, for x-oriented electric field across $e]_{c}^{90}$, the $e_{33}$ now is at 1,1 position which couples x-field to x -longitudinal stress, and $e_{31}$ at the 1,3 position which couples x-field to z- transverse stresses. This matches the elliptical motion that builds the Rayleigh wave.
Fig. S5: Simulated acoustic mode profiles. Simulated mode profiles with its corresponding driving domain for (a) Rayleigh wave and (b) Bulk shear wave.
1.5 Single domain piezoelectric coefficient of twin-domain structure
Starting from the electromechanical coupling coefficient definition (?)
$K=\frac{U_{m}}{\sqrt{U_{e}U_{d}}},$ (S13)
where $U_{m}$ is the mutual energy, $U_{e}$ is the elastic energy and $U_{d}$ is the dielectric energy. For a twin-domain SrTiO3 thin film with $a$- and $c$-oriented domains, the energies for an electric field oriented along the crystallographic $c$ axis are given by
$U_{m}=\frac{1}{2}\left[\int_{\Omega_{a}}\left(\frac{e_{33}^{(a)}}{c_{33}^{E(a)}}\right)E_{3}T_{3}dV+\int_{\Omega_{c}}\left(\frac{e_{33}^{(c)}}{c_{33}^{E(c)}}\right)E_{3}T_{3}dV\right],$ (S14)
$U_{e}$ $=\frac{1}{2}\int_{\Omega}\frac{1}{c_{33}^{E}}T_{3}^{2}dV,$ (S15)
$U_{d}$ $=\frac{1}{2}\int_{\Omega}\varepsilon_{33}^{T}E_{3}^{2}dV.$ (S16)
Here, $E_{3}$ and $T_{3}$ are the electric field and the stress in the c-axis, respectively. For our twin boundary, the volumetric fraction between $\mathbf{(c)}$-and $\mathbf{(a)}$-oriented domains is 50% in the case of 0 and 90 degree IDTs
$\operatorname{Vol}\left(\Omega_{a}\right)=\operatorname{Vol}\left(\Omega_{c}\right)=\frac{1}{2}\operatorname{Vol}(\Omega).$ (S17)
Since a-oriented domains do not excite Rayleigh waves ($e_{33}^{(a)}=0$), only the $c$-domain volume contributes to the mutual energy, while the elastic and dielectric energies are calculated over both c and a volume/domain. As a result, the electromechanical coupling coefficient that was utilized in the acousto-optic modulator was reduced by a factor of two relative to a single-domain $c$-oriented film
$K_{33,\mathrm{eff}}=\frac{1}{2}K_{33}^{(c)}.$ (S18)
Equivalently, the effective piezoelectric coefficient satisfies $e_{33,\mathrm{eff}}=e_{33}^{(c)}/2$.
Similarly, for the $e_{15}$ with a 45 IDT, the mutual energy is given by
$U_{m}=\frac{1}{2}\left[\int_{\Omega_{x}}\left(\frac{e_{15}^{(x)}}{c_{55}^{E(x)}}\right)\cos(45^{\circ})^{2}E_{1}T_{5}dV+\int_{\Omega_{y}}\left(\frac{e_{15}^{(y)}}{c_{55}^{E(y)}}\right)\cos(45^{\circ})^{2}E_{1}T_{5}dV\right].$ (S19)
where the x, and y are the in-plane orientation 0,and 90. The total energies are,
$U_{e}$ $=\frac{1}{2}\int_{\Omega}\frac{1}{c_{55}^{E}}T_{5}^{2}dV,$ (S20)
$U_{d}$ $=\frac{1}{2}\int_{\Omega}\varepsilon_{11}^{T}E_{1}^{2}dV.$ (S21)
Since both directions exhibit identical response ($e_{15}^{(x)}=e_{15}^{(y)}=e_{15}^{(a)}$), while the volumetric fraction of the domains is still 50%,
$K_{15,\mathrm{eff}}=\frac{1}{2}K_{15}^{(a)}.$ (S22)
Equivalently, the effective piezoelectric coefficient satisfies $e_{15,\mathrm{eff}}=e_{15}^{(a)}/2$.
1.6 Acoustic propagation losses
A propagating acoustic mode in a multilayer structure can be described by a complex wavevector
$\tilde{k}=\beta+i\alpha,$ (S23)
where $\beta$ is the real propagation constant and $\alpha>0$ quantifies spatial attenuation. The displacement field therefore decays as
$u(x)=u_{0}e^{i\beta x}e^{-\alpha x}.$ (S24)
Since the strain field $\varepsilon(x)$ is proportional to the spatial derivative of the displacement, $\varepsilon\sim\partial u/\partial x$, its amplitude inherits the same exponential decay,
$\varepsilon(x)\propto e^{-\alpha x}.$ (S25)
We thus define the strain amplitude decay length as
$L_{A}\equiv\frac{1}{\alpha},$ (S26)
corresponding to the propagation distance over which the strain amplitude decays to $1/e$ of its initial value. The quality factor $Q$ of the propagating mode can be expressed directly in terms of the real and imaginary parts of the wavevector as
$Q=\frac{\beta}{2\alpha},$ (S27)
[…]
which reflects the ratio between phase accumulation and amplitude loss per propagation cycle. Solving for $\alpha$ and substituting into the definition of $L_{A}$ yields
$L_{A}=\frac{2Q}{\beta}=\frac{Q\,\lambda}{\pi}.$ (S28)
In our device shown in Fig. S6 ($f_{0}=504~\mathrm{MHz}$, $\gamma/2\pi=1.136~\mathrm{MHz}$), the loaded quality factor is $Q_{L}=f_{0}/(\gamma/2\pi)\approx 444$. With an external coupling fraction $\gamma_{\mathrm{ext}}/\gamma=0.1447$, we obtain $Q_{i}=Q_{L}/(1-\gamma_{\mathrm{ext}}/\gamma)\approx 519$. For an acoustic wavelength $\lambda=8~\mu\mathrm{m}$, the intrinsic strain decay length is
$L_{A,i}=\frac{Q_{i}\lambda}{\pi}\approx\frac{519\times 8~\mu\mathrm{m}}{\pi}\approx 1.32~\mathrm{mm}.$ (S29)
2 Acousto-optic modulation
2.1 Calculating the $V_{\pi}L$
To accurately calculate the $V_{\pi}L$ of the acousto-optic modulator, we measure the electrical output coming from the photodetector using an oscilloscope as shown in Fig. S7. We modulate the ring resonator at the 3-dB point to maximize the modulation depth while driving the RF at the Rayleigh wave resonance frequency with 1 dBm RF power. Using the ring transfer function, we can relate the photodetector output peak to peak voltage ($V_{pp}$) to the wavelength shift ($\Delta\lambda$) caused by the $\Delta n_{\mathrm{eff}}$.
Refer to caption(a)
Refer to caption(b)
Fig. S7: Acousto-optic modulation measurement. (a) Racetrack response showing the 8.4 pm wavelength shift. (b) Averaged oscilloscope output voltage.
Since the photodetector voltage is linear with the optical power, a change of 2.21 dB in the output voltage corresponds to an identical 2.21 dB change in the optical power. This power variation corresponds to a wavelength shift of $\Delta\lambda=8.4~\mathrm{pm}$.
The power and thus the driving voltage delivered to the device can be calculated from the $S_{11}$ measurement as
$P_{\mathrm{device}}=P_{\mathrm{input}}(1-|S_{11}|^{2}),$ (S30)
where $P_{\mathrm{input}}$ is the input RF power of the VNA. The peak voltage is defined as $V_{p}=\sqrt{2\;Z_{L}\;P_{\mathrm{device}}}$ where $Z_{L}$ is the IDT impedance. The calculated peak voltage is $V_{p}=0.235\;V$ for 1 dBm input power. The $V_{\pi}L$ can be calculated as
$V_{\pi}L=\frac{FSR\;V_{p}\;L}{\Delta\lambda},$ (S31)
where $FSR$ denotes the free spectral range of the ring resonator, $FSR=0.781~\mathrm{nm}$, and $L_{c}$ is the interaction length between the optical mode and the acoustic wave. Using the above equation, we extract $V_{\pi}L=0.874\pm 0.084~\mathrm{V\cdot cm}$ for an interaction length of $L=2\times 200~\mu\mathrm{m}$.
[…]
2.2 Photoelasticity
The aforementioned wavelength shift, $\Delta\lambda$, arises from a change in the effective refractive index $n_{\mathrm{eff}}$ of the ring resonator. The relationship between $\Delta\lambda$ and $\delta n_{\mathrm{eff}}$ can be obtained by writing the resonance phase condition for the $m^{\text{th}}$ azimuthal mode in the resonator
$\beta^{(0)}\left(\omega_{0}\right)L=2\pi m,$ (S32)
while the perturbed state satisfies
$\beta^{(1)}\left(\omega_{1}\right)L_{c}+\beta^{(0)}\left(\omega_{1}\right)\left(L-L_{c}\right)=2\pi m,$ (S33)
where $L$ is the circumference of the racetrack, given by its radius $R$ and straight length $L_{c}/2$ (i.e. L= $2\pi R+L_{c}$), and $\beta^{(i)}(\omega_{j})$ is the propagation constant of the waveguide mode ($i=0$ non-perturbed, $i=1$ perturbed) evaluated at a specific frequency. Subtracting both equations we get
$\left(\beta^{(1)}\left(\omega_{1}\right)-\beta^{(0)}\left(\omega_{1}\right)\right)L_{c}+\left(\beta^{(0)}\left(\omega_{1}\right)-\beta^{(0)}\left(\omega_{0}\right)\right)L=$  $0$ (S34)
$\left(\frac{\omega_{0}}{c}\delta n_{eff}\right)L_{c}+\left(\frac{n_{g}}{c}\delta\omega\right)L=$  $0,$
we obtain
$\delta n_{\mathrm{eff}}=-n_{g}\frac{\delta\omega}{\omega}\frac{L}{L_{c}},$ (S35)
where $n_{g}$ is the group index and $\omega$ is the mode frequency. The term $\delta\omega/\omega$ can be cast into a wavelength form with a change of variables ($\delta\omega/\omega=-\delta\lambda/\lambda$), while $n_{g}$ can be inferred from the $\mathrm{FSR}$ via $n_{g}=\lambda^{2}/(\mathrm{FSR}\times L)$. resulting in the final expression for the measured $\delta n_{eff}$ as
$\delta n_{\mathrm{eff}}=\frac{\lambda\;\delta\lambda}{FSR\times L_{c}}.$ (S36)
Hence, the total effective refractive index modulation can be directly obtained from the measurement.
$\delta n_{\mathrm{eff}}$ can also be inferred from perturbation theory (?)
$\delta n_{\mathrm{eff}}=\frac{n_{\mathrm{eff}}}{2}\frac{\left\langle E\right|\Delta\varepsilon\left|E\right\rangle}{\left\langle E\right|\varepsilon\left|E\right\rangle},$ (S37)
where $\Delta\varepsilon$ is the permittivity perturbation and $\left|E\right\rangle$ is the unperturbed field. The effective refractive index modulation $\delta n_{\mathrm{eff}}$ consists of contributions arising from moving boundaries $(\delta n_{\mathrm{eff,MB}})$ and from the photoelastic effect $(\delta n_{\mathrm{eff,PE}})$. The photoelastic contribution is further decomposed into the contributions originating from the CSAR waveguide $(\delta n_{\mathrm{eff,CSAR}})$, the $\mathrm{SiO}_{2}$ substrate $(\delta n_{\mathrm{eff,SiO_{2}}})$, and the SrTiO3 layer $(\delta n_{\mathrm{eff,STO}})$.
To isolate $\delta n_{\mathrm{eff,STO}}$, all other contributions are calculated independently using finite‑element simulations. A surface acoustic wave is simulated in the cross‑section of the ring resonator to obtain the strain and displacement fields at the waveguide, while a separate electromagnetic eigenmode simulation provides the unperturbed optical field. The individual contributions are subsequently evaluated by post‑processing the simulated fields. In this simulation, the extracted effective piezoelectric and dielectric constant of the experiment are used at $3~V/\mu m$ ($e_{33,\mathrm{eff}}=5.52\;\mathrm{C/m}^{2}$ & $\epsilon=658$).
2.2.1 Moving boundary contribution
The perturbation term in Eq. (S37) arising from domains with shifting material boundaries was first derived in Ref.  (?). For an interface between two materials with permittivities $\varepsilon_{1}$ and $\varepsilon_{2}$, displaced by a mechanical displacement field $\mathbf{u}$, the corresponding moving‑boundary contribution to the permittivity perturbation can be expressed as
$\left\langle E\right|\Delta\varepsilon\left|E\right\rangle=\int ds\ \hat{n}\cdot\vec{u}\left[(\varepsilon_{1}-\varepsilon_{2})\left|E_{\|}\right|^{2}-(\varepsilon_{1}^{-1}-\varepsilon_{2}^{-1})\left|D_{\perp}\right|^{2}\right],$ (S38)
where $\hat{n}$ denotes the unit normal to the material interface, $E_{\parallel}$ is the component of the electric field parallel to the interface, and $D_{\perp}$ is the component of the electric displacement field normal to the interface. The surface integral is evaluated over the material boundary stated as $ds$. For the waveguide considered here, the total moving‑boundary contribution is obtained by summing the contributions from all material interfaces in the waveguide cross‑section.
2.2.2 Photoelastic contribution
The photoelastic contribution $\delta n_{\mathrm{eff,PE},\mathrm{M}}$ of each material (M) besides SrTiO3, namely CSAR and SiO2 is given by
$\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{M}}=\frac{\epsilon_{0}n_{\mathrm{M}}^{4}n_{\mathrm{eff}}}{2}\frac{\int dr\left(\begin{array}[]{lll}E_{x}^{*}&E_{y}^{*}&E_{z}^{*}\end{array}\right)\left(\begin{array}[]{lll}dB_{1}&dB_{6}&dB_{5}\\ dB_{6}&dB_{2}&dB_{4}\\ dB_{5}&dB_{4}&dB_{3}\end{array}\right)\left(\begin{array}[]{l}E_{x}\\ E_{y}\\ E_{z}\end{array}\right)}{\int E_{\;}^{\ast}\epsilon E\,dr},$ (S39)
where $dB_{ij}=\Delta(1/\varepsilon_{ij})$ denotes the change in the optical impermeability tensor induced by the photoelastic effect, written here in Voigt notation. We consider that for all materials in the waveguide, the photoelastic effect equation can be written as:
$\left(\begin{array}[]{c}dB_{1}\\ dB_{2}\\ dB_{3}\\ dB_{4}\\ dB_{5}\\ dB_{6}\end{array}\right)=\left(\begin{array}[]{cccccc}p_{11}&p_{12}&p_{12}&0&0&0\\ p_{12}&p_{11}&p_{12}&0&0&0\\ p_{12}&p_{12}&p_{11}&0&0&0\\ 0&0&0&p_{44}&0&0\\ 0&0&0&0&p_{44}&0\\ 0&0&0&0&0&p_{44}\end{array}\right)\left(\begin{array}[]{c}s_{1}\\ s_{2}\\ s_{3}\\ s_{4}\\ s_{5}\\ s_{6}\end{array}\right),$ (S40)
where $s$ is the strain in the material in the Voigt notation. For the wave considered here, the only non‑zero strain components are $s_{11}\,(s_{1})$, $s_{33}\,(s_{3})$, and $s_{13}\,(s_{5})$, resulting in the following non‑zero components of the impermeability tensor
$dB_{1}$ $=p_{11}s_{1}+p_{13}s_{3},$ (S41)
$dB_{3}$ $=p_{31}s_{1}+p_{33}s_{3},$
$dB_{5}$ $=p_{44}s_{5}.$
With this, we rewrite the photoelastic contribution:
$\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{M}}=-\frac{\epsilon_{0}n_{\mathrm{M}}^{4}n_{\mathrm{eff}}}{2}\frac{\int dr\ [E_{x}^{*}dB_{1}E_{x}+E_{z}^{*}dB_{3}E_{z}+dB_{5}(E_{x}^{*}E_{z}+E_{x}E_{z}^{*})]}{\int E^{*}\epsilon Edr},$ (S42)
which gives us the photoelastic contributions of $\mathrm{CSAR}$ and $\mathrm{SiO}_{2}$. In the absence of reported photoelastic coefficients for CSAR 62, we approximate its stress–optic response using those of PMMA (?), a closely related polymeric e-beam resist, noting that the opto-acoustic coupling is dominated by the waveguide core and that the resist contributes only through a weak evanescent optical overlap ($28\%$).
2.2.3 Effective photoelastic coefficient of SrTiO3
As the individual tensor components of the photoelastic response of SrTiO3 cannot be independently resolved, we introduce an effective photoelastic coefficient $p_{\mathrm{eff}}$. By setting all non‑zero elements of the photoelastic tensor equal to $p_{\mathrm{eff}}$, the photoelastic contribution can be expressed as
$\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{STO}}$ $=\frac{\epsilon_{0}n_{\mathrm{STO}}^{4}n_{\mathrm{eff}}}{2}\frac{\int dr\ [(s_{1}+s_{3})(|E_{x}|^{2}+|E_{z}|^{2})+2s_{5}\mathrm{Re}(E_{x}^{*}E_{z})]}{\int E^{*}\epsilon Edr}p_{\mathrm{eff}},$ (S43)
$=\mathcal{L}\;p_{\mathrm{eff}}.$
The constant $\mathcal{L}$ can thus be extracted by post‑processing the simulated strain and optical mode profiles, and subsequently used to determine the effective photoelastic coefficient $p_{\mathrm{eff}}$ as:
$p_{\mathrm{eff}}=\frac{1}{\mathcal{L}}(\delta n_{\mathrm{eff}}-\delta n_{\mathrm{eff},\mathrm{MB}}-\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{CSAR}}-\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{SiO}_{2}}).$ (S44)
Table S1 summarizes the analysis parameters and results. Finally, we benchmark the extracted effective photoelastic coefficient of SrTiO3 against that of BaTiO3, using identical simulated strain and optical field distributions. The photoelastic tensor elements for BaTiO3 are taken from literature  (?)
$p_{\mathrm{eff,BTO}}=\frac{\int dr\left[E_{x}^{*}dB_{1,BTO}E_{x}+E_{y}^{*}dB_{2,BTO}E_{y}+E_{z}^{*}dB_{3,BTO}E_{z}+dB_{5,BTO}\left(E_{x}^{*}E_{z}+E_{x}E_{z}^{*}\right)\right]}{\int dr\left[\left(s_{1}+s_{3}\right)\left(\left|E_{x}\right|^{2}+\left|E_{y}\right|^{2}+\left|E_{z}\right|^{2}\right)+2s_{5}\operatorname{Re}\left(E_{x}^{*}E_{z}\right)\right]},$ (S45)
yielding $p_{\mathrm{eff,BTO}}=0.61$. This direct comparison highlights the relative strength of the photoelastic response in SrTiO3.
Table S1: Key parameters and extracted quantities used in this work.
Parameter Value
$R$ $200\,\mu\mathrm{m}$
$L_{c}$ $400\,\mu\mathrm{m}$
$\lambda$ $1550.1\,\mathrm{nm}$
FSR $0.781\,\mathrm{nm}$
$n_{g}$ $1.857$
$\delta n_{\mathrm{eff}}$ $2.0840\times 10^{-5}$
$\delta n_{\mathrm{eff,MB}}$ $1.263\times 10^{-6}$
$\delta n_{\mathrm{eff,PE(CSAR+SiO_{2})}}$ $4.754\times 10^{-6}$
$\mathcal{L}$ $2.635\times 10^{-5}$
$p_{\mathrm{eff,STO}}$ $0.5625$
2.3 Optical response with increased RF power
Fig. S8: Optical response vs. RF power. (a) Measured optical transmission spectra of the ring resonator for different RF powers applied to the IDT. (b) Simulated optical transmission.
Fig. S8a shows the optical transmission as a function of the applied RF power. In the unresolved-sideband regime $\Omega\ll\kappa$, the optical cavity responds adiabatically to the mechanical motion, such that the optomechanical interaction induces resonance-frequency jitter of the cavity. The measured transmission, which represents a time average over this frequency jitter, therefore exhibits an apparent broadening relative to the intrinsic cavity linewidth. Additionally, we observed that increasing the RF power leads to a slight increase in the sample temperature, accompanied by a blue shift of the optical resonance. Using temperature-dependent measurements of the ring response, we correlated the observed thermal shift with the actual device temperature, estimating a temperature increase of 6.26 K at 10 dBm RF power.
The broadening of the measured optical response under high RF power can be numerically modeled assuming a weak optical input using the equation of the intracavity field amplitude $a(t)$ (?)
$\dot{a}(t)=\left(i\Delta-\frac{\kappa}{2}\right)a(t)-i\,\delta\omega\cos(\Omega t)\,a(t)+\sqrt{\kappa_{e}}\,a_{\mathrm{in}},$ (S46)
where $\Delta=\omega_{L}-\omega_{c}$ is the laser–cavity detuning, $\kappa=\kappa_{e}+\kappa_{i}$ is the total optical decay rate, and $\delta\omega$ denotes the amplitude of the acoustically induced modulation of the cavity resonance frequency. For a periodic modulation, the steady-state solution is expanded as
$a(t)=\sum_{n}a_{n}e^{-in\Omega t},$ (S47)
yielding the coupled equations
$\left(\frac{\kappa}{2}-i(\Delta+n\Omega)\right)a_{n}+i\frac{\delta\omega}{2}\left(a_{n-1}+a_{n+1}\right)=\sqrt{\kappa_{e}}\,a_{\mathrm{in}}\,\delta_{n0}.$ (S48)
The through-port output field is
$a_{\mathrm{out}}(t)=a_{\mathrm{in}}-\sqrt{\kappa_{e}}\,a(t),$ (S49)
because the optical transmission is recorded using a low-bandwidth optical power meter, whose response time is much longer than the modulation period $2\pi/\Omega$, the measured signal corresponds to the time-averaged transmission
$T_{\mathrm{DC}}=\left\langle\frac{|a_{\mathrm{out}}(t)|^{2}}{|a_{\mathrm{in}}|^{2}}\right\rangle=\frac{\left|a_{\mathrm{in}}-\sqrt{\kappa_{e}}\,a_{0}\right|^{2}+\kappa_{e}\sum_{n\neq 0}|a_{n}|^{2}}{|a_{\mathrm{in}}|^{2}}.$ (S50)
The quasi-DC transmission spectrum $T_{\mathrm{DC}}(\lambda)$ is then obtained by evaluating Eq. (S50) as a function of laser detuning $\Delta(\lambda)$, as shown in Fig. S8b.
3 Materials overview
The performance of piezo-optomechanical transducers is ultimately dictated by the material on which they are realized. Material properties determine the strength of both the electromechanical and optomechanical interactions that underpin coherent microwave-optical transduction. In this section, we compare several widely used piezoelectric material with the aim of assessing their suitability for efficient piezo-optomechanical devices.
Fig. S9: Piezoelectric tensor comparison. Piezoelectric tensor elements of commonly used materials. Effective parameters in SrTiO3, arising from twin domain boundaries, are denoted by stars. (?, ?, ?, ?, ?, ?).
From an electromechanical perspective, a central figure of merit is the electromechanical coupling rate $g_{\mathrm{EM}}$. In piezoelectric resonators, $g_{\mathrm{EM}}$ scales with the square root of the effective electromechanical coupling coefficient $k^{2}$ (?), which is determined by the relevant piezoelectric tensor elements $e_{ij}$, as well as by device geometry and modal overlap. Materials with large intrinsic piezoelectric coefficients, such as AlN and LiNbO3, naturally support large $k^{2}$ and can therefore achieve strong electromechanical coupling in compact geometries. In contrast, GaAs exhibits a comparatively weak piezoelectric response, resulting in smaller $k^{2}$ and correspondingly reduced $g_{\mathrm{EM}}$, unless additional resonant enhancement or impedance-matching strategies are employed. Fig. S9 compares the dominant piezoelectric tensor elements for the materials considered here.
At cryogenic temperatures, the piezoelectric response of ferroelectric materials generally decreases as the operating temperature moves further away from the Curie point. The magnitude of this reduction is material dependent; for example, LiNbO3 exhibits a relatively modest decrease of approximately 6% at 1.3 K (?).
Optomechanical performance, by contrast, is primarily characterized by the single-photon optomechanical coupling rate $g_{0}$, which depends on the photoelastic tensor, the refractive index, and the achievable optical mode confinement. While conventional piezoelectric materials provide moderate photoelastic coupling, hybrid photonic platforms enable access to materials with substantially larger photoelastic coefficients. In particular, SiN-on-SrTiO3 and all-SrTiO3 waveguides would leverage a strong photoelastic response enabling enhanced optomechanical interaction strengths. Table S2 summarizes the relevant photoelastic tensor elements for representative optical guiding materials.
Table S2: Photoelastic tensor elements of commonly used optical guiding materials, where bold numbers indicate the high values (?, ?, ?, ?).
Material Selected photoelastic tensor elements
Si $\mathbf{p_{11}=-0.094},\;p_{12}=0.017,\;p_{44}=-0.051$
SiN $\mathbf{\left|p_{44}\right|\approx 0.086},p_{11}=0.047$
LiNbO3 $\mathbf{p_{31}=0.179},\;p_{33}=0.071,\;p_{44}=0.146,\;p_{41}=-0.151,\dots$
BaTiO3 $\mathbf{p_{44}=1.00},\;p_{11}=0.50,\;p_{33}=0.77,\;p_{12}=0.106,\;p_{13}=0.20,\dots$
SrTiO3 $p_{\mathrm{eff}}\approx 0.5625$